# Joint Velocity-Growth Flow Matching for Single-Cell Dynamics Modeling

### v1

@@ B.3 Dyngen Data
Dyngen is a multi-modal simulation engine for studying dynamic cellular processes at single-cell resolution [52]. We follow the setup of [30], which simulates gene expression time-series data that mimics cell proliferation processes, including branching and temporal progression. In the analyzed instance of the dyngen dataset, the samples span five discrete time points, labeled from 0 to 4, capturing the temporal evolution of the cell population.
We use PHATE [53] to reduce its dimensions to 5, the same as [30]. Starting from time point 0, cells can be divided into two branches based on the sign of the second PHATE coordinate ($x_{2}$). Quantitative analysis reveals a pronounced branch imbalance: for instance, at time point 4, 88 cells belong to one branch while 213 belong to the other, which poses a great challenge to the models that do not consider unbalancedness, resulting in cross-branch inference. Throughout this study, we assume that the temporal change in cell counts within each branch is entirely governed by a growth function, without contributions from migration or observational noise.
By setting $\tau=5$ and $\epsilon=0.03$ according to our selection scheme described in Appendix B.1 and illustrated in Fig. LABEL:fig:tau (b), we train the model with a warm-up stage of 400 iterations. After that, distribution fitting loss $\mathcal{L}_{\rm OT}$ is introduced for an additional 30 training epochs. Our model successfully avoids cross-branch reconstructions in trajectory modeling, as illustrated in Fig. LABEL:fig:dygen_vis. This leads to generated samples that remain well-aligned with the underlying manifold. In terms of growth rate estimation, our model also accurately captures the rapid expansion observed in the branch below between time point 3 and time point 4, where the number of observed cells increases dramatically from 25 to 213, indicating the highest growth rate in the system.
@@ B.4 Gausssian 1000D Data
DeepRUOT [25] employs a high-dimensional Gaussian mixture distribution (100D) [54] to evaluate the scalability of models. In this work, we adopt an even more challenging setting by testing scalability at 1000 dimensions. Following the setup of [25], for the initial distribution, we generated 400 samples from the Gaussian located lower in the $(x_{1},x_{2})$ plane, and 100 samples from the Gaussian positioned higher. For the final distribution, we generated 1,000 samples from the upper Gaussian, and 200 samples each from the two lower Gaussians, and assume cells in the upper region exhibit proliferation without transport [5]. We observe that simulation-based methods [25] encounter training instability in the 1000-dimensional setting and flow matching-based methods fail to learn a reliable velocity field due to the unbalancedness of data, as shown in Tab. 1. (Some works incorporating unbalancedness [35, 40] may as well reconstruct correct velocity field but fail to construct growth rate). In contrast, By setting $\tau=5$ and $\epsilon=0.03$ according to our selection scheme described in Appendix B.1 and illustrated in Fig. LABEL:fig:tau (c), we train the model with a warm-up stage of 500 iterations. After that, distribution fitting loss $\mathcal{L}_{\rm OT}$ is introduced for an additional 10 training epochs. Our method remains effective even in high-dimensional regimes for both velocity and growth field learning.
@@ B.5 Additional Experiment: Mouse Hematopoiesis Data
We also validate our algorithm on the mouse hematopoiesis data previously analyzed in [60, 22, 25]. This dataset leverages lineage tracing to track differentiation trajectories. After applying batch correction to integrate data across multiple experiments, the cells were embedded into a two-dimensional force-directed layout (SPRING plot). The resulting visualization reveals a pronounced bifurcation, where early progenitor cells diverge into two distinct differentiation lineages.
Our model effectively captures the underlying branching structure in the data, and the predicted growth rates align well with those reported in [25], as well as with established biological priors and are consistent with known biological lineage patterns. Following the evaluation protocol of [25], we also compute the Wasserstein-1 distance as a quantitative metric. Under this evaluation, by setting $\tau=20$ and $\epsilon=0.005$ according to our selection scheme described in Appendix B.1, we train the model with a warm-up stage of 200 iterations. After that, the distribution fitting loss $\mathcal{L}_{\rm OT}$ is introduced for an additional 50 training epochs. Our method achieves superior performance, demonstrating improved trajectory inference. In addition, we evaluate the RME (Relative Mass Error) metric introduced in this paper. The results indicate that our model significantly outperforms the other baselines in terms of mass-matching reconstruction accuracy, as shown in Tab. A-6 and Fig. A-16.
Table A-6: $\mathcal{W}_{1}$ and RME over all time on mouse hematopoiesis data. Part of the results were adopted from [25].
@@
SF2[39] 0.167 0.190
uAM[42] 0.745 0.777
UDSB [49] 0.388 0.159 0.128 0.249
TIGON [22] 0.314 0.124 0.342 0.177
DeepRUOT [25] 0.145 0.140 0.132 0.202
@@ B.6 EB Data
Extensive experiments on the Embryoid Body (EB) [55] dataset with varying PCA dimensions are conducted in Sect. 4. We preprocess this dataset first by normalizing EB (5D) and leaving EB (50D) unnormalized. The proposed VFGM demonstrates strong performance in both cellular dynamics reconstruction and growth prediction, as illustrated in Fig. LABEL:fig:eb_vis. Specifically, we set the parameters as $\epsilon=0.01$, $\tau=30$, $M_{1}=800$, and $M_{2}=830$, and normalize the cost matrix to ensure numerical stability when computing the semi-relaxed optimal transport plans (Eq. (11)).
@@
Big batch strategy for large-scale datasets. When the number of observed data points is large, computing the transport plan between two consecutive time points using Eq. (11) becomes computationally inefficient. To address this, we adopt a Big-Batch Strategy. Specifically, we partition the data at each time point into $n$ subsets, referred to as Big Batches. For each of the $n$ Big Batches, we precompute and store the transport plans between adjacent time points. During training, we randomly select one Big Batch and sample smaller mini-batches from it based on the precomputed transport plan for model optimization. In real-world datasets, we set $n=5$, while in synthetic datasets above we do not apply this strategy, i.e., $n=1$, as the sample size is relatively smaller. The mini-batch size is set to 256 for all experiments, except for Dyngen, where we use a smaller batch size of 32 due to its limited number of samples.
@@
We also examine whether the distribution fitting loss might lead to overfitting on the given population snapshots. In the hold-one-out experiments (Tab. 2), we relabel the timestamps of the four timepoints as 0, 1, 2, 3, and 4. Following the protocol in [36], we evaluate the model by holding out one intermediate timepoint at a time (i.e., 1, 2, and 3) and report the average Wasserstein-1 distance. While hold-one-out experiments have already validated the superiority of VFGM over other methods, we further analyze the loss curves of both the distribution fitting loss and the $\mathcal{W}_{1}$ distance on a hold-out distribution (e.g., the distribution at time 1) that is not seen during training (Fig. A-18). The results show that the $\mathcal{W}_{1}$ distance on the hold-out distribution decreases during the optimization of $\mathcal{L}_{\rm OT}$, indicating the absence of overfitting and demonstrating the generalization ability of the distribution fitting loss in modeling unseen distributions.
@@ B.7 CITE Data
CITE-seq (Cellular Indexing of Transcriptomes and Epitopes by Sequencing) [56] is an advanced technique that enables the simultaneous profiling of transcriptomes and surface protein expression at the single-cell level through the use of antibody-derived tags. In this study, we utilize only the gene expression matrix from the CITE-seq dataset, and preprocess the data by normalizing CITE (5D) and leaving CITE (50D) unnormalized. The experimental settings for the CITE dataset are consistent with those used in the EB data experiments, employing PCA with 5 and 50 dimensions, and hyperparameters set as $\epsilon=0.01$, $\tau=30$, $M_{1}=800$, and $M_{2}=830$. We relabel the timestamps of the four timepoints as 0, 1, 2, and 3. Following the protocol in [36], we evaluate the model by holding out one intermediate timepoint at a time (i.e., 1 and 2) and report the average Wasserstein-1 distance.
@@
To assess the model’s capability to capture state trajectories (Fig. LABEL:fig:cite_vis (a)) and predict mass growth (Fig. LABEL:fig:cite_vis (b)), the distribution at time point 1 is held out. Additionally, the distribution at time point 2 is held out to evaluate the performance of the $\mathcal{L}_{\rm OT}$ and $\mathcal{W}_{1}$ distance (hold-out) curves (Fig. LABEL:fig:cite_hoo). The strong performance observed in the visualizations substantiates the potential of the proposed VGFM framework.

### v2

@@ B.3 Dyngen Data
Dyngen is a multi-modal simulation engine for studying dynamic cellular processes at single-cell resolution [cannoodt2021spearheading]. We follow the setup of [huguet2022manifold], which simulates gene expression time-series data that mimics cell proliferation processes, including branching and temporal progression. In the analyzed instance of the dyngen dataset, the samples span five discrete time points, labeled from 0 to 4, capturing the temporal evolution of the cell population.
We use PHATE [moon2019visualizing] to reduce its dimensions to 5, the same as [huguet2022manifold]. Starting from time point 0, cells can be divided into two branches based on the sign of the second PHATE coordinate ($x_{2}$). Quantitative analysis reveals a pronounced branch imbalance: for instance, at time point 4, 88 cells belong to one branch while 213 belong to the other, which poses a great challenge to the models that do not consider unbalancedness, resulting in cross-branch inference. Throughout this study, we assume that the temporal change in cell counts within each branch is entirely governed by a growth function, without contributions from migration or observational noise.
By setting $\tau=5$ and $\epsilon=0.03$ according to our selection scheme described in Appendix B.1 and illustrated in Fig. 14(a) (b). Our model successfully avoids cross-branch reconstructions in trajectory modeling, as illustrated in Fig. 16(a). This leads to generated samples that remain well-aligned with the underlying manifold. In terms of growth rate estimation, our model also accurately captures the rapid expansion observed in the branch below between time point 3 and time point 4, where the number of observed cells increases dramatically from 25 to 213, indicating the highest growth rate in the system.
(a) Predicted dynamics and growth rates by VGFM on Dyngen data.
@@ B.4 Gausssian 1000D Data
DeepRUOT [zhang2024learning] employs a high-dimensional Gaussian mixture distribution (100D) [ruthotto2020machine] to evaluate the scalability of models. In this work, we adopt an even more challenging setting by testing scalability at 1000 dimensions. Following the setup of [zhang2024learning], for the initial distribution, we generated 400 samples from the Gaussian located lower in the $(x_{1},x_{2})$ plane, and 100 samples from the Gaussian positioned higher. For the final distribution, we generated 1,000 samples from the upper Gaussian, and 200 samples each from the two lower Gaussians, and assume cells in the upper region exhibit proliferation without transport [zhang2024learning]. We observe that simulation-based methods [zhang2024learning] encounter training instability in the 1000-dimensional setting and flow matching-based methods fail to learn a reliable velocity field due to the unbalancedness of data, as shown in Tab. 1. (Some works incorporating unbalancedness [eyringunbalancedness, corso2025composing] may as well reconstruct correct velocity field but fail to construct growth rate). In contrast, By setting $\tau=5$ and $\epsilon=0.03$ according to our selection scheme described in Appendix B.1 and illustrated in Fig. 14(a) (c). Our method remains effective even in high-dimensional regimes for both velocity and growth field learning.
(b) Predicted dynamics and growth rates by VGFM on Gaussian 1000D data.
@@ B.5 Additional Experiment: Mouse Hematopoiesis Data
We also validate our algorithm on the mouse hematopoiesis data previously analyzed in [weinreb2020lineage, sha2024reconstructing, zhang2024learning]. This dataset leverages lineage tracing to track differentiation trajectories. After applying batch correction to integrate data across multiple experiments, the cells were embedded into a two-dimensional force-directed layout (SPRING plot). The resulting visualization reveals a pronounced bifurcation, where early progenitor cells diverge into two distinct differentiation lineages.
(c) Predicted dynamics and growth rates by VGFM on mouse hematopoiesis data.
Our model effectively captures the underlying branching structure in the data, and the predicted growth rates align well with those reported in [zhang2024learning], as well as with established biological priors and are consistent with known biological lineage patterns. Following the evaluation protocol of [zhang2024learning], we also compute the Wasserstein-1 distance as a quantitative metric. Under this evaluation, by setting $\tau=20$ and $\epsilon=0.005$ according to our selection scheme described in Appendix B.1, we train the model with a warm-up stage of 200 iterations. After that, the distribution fitting loss $\mathcal{L}_{\rm OT}$ is introduced for an additional 100 training epochs. Our method achieves superior performance, demonstrating improved trajectory inference. In addition, we evaluate the RME (Relative Mass Error) metric introduced in this paper. The results indicate that our model significantly outperforms the other baselines in terms of mass-matching reconstruction accuracy, as shown in Tab. A-9 and Fig. A-16.
Table A-9: $\mathcal{W}_{1}$ and RME over all time on mouse hematopoiesis data. Part of the results were adopted from [zhang2024learning].
@@
SF2[tong2024simulation] 0.167 0.190
uAM[neklyudov2023action] 0.745 0.777
UDSB [pariset2023unbalanced] 0.388 0.159 0.128 0.249
TIGON [sha2024reconstructing] 0.314 0.124 0.342 0.177
DeepRUOT [zhang2024learning] 0.145 0.140 0.132 0.202
@@ B.6 EB Data
Extensive experiments on the Embryoid Body (EB) [moon2018manifold] dataset with varying PCA dimensions are conducted in Sect. 4. We preprocess this dataset first by normalizing EB (5D) and leaving EB (50D) unnormalized. The proposed VFGM demonstrates strong performance in both cellular dynamics reconstruction and growth prediction, as illustrated in Fig. 18(a). Specifically, we set the parameters as $\epsilon=0.01$, $\tau=5$ and normalize the cost matrix to ensure numerical stability when computing the semi-relaxed optimal transport plans (Eq. (11)).
@@
Big batch strategy for large-scale datasets. When the number of observed data points is large, computing the transport plan between two consecutive time points using Eq. (11) becomes computationally inefficient. To address this, we adopt a Big-Batch Strategy. Specifically, we partition the data at each time point into $n$ subsets, referred to as Big Batches. For each of the $n$ Big Batches, we precompute and store the transport plans between adjacent time points. During training, we randomly select one Big Batch and sample smaller mini-batches from it based on the precomputed transport plan for model optimization. In real-world datasets, we set $n=5$, while in synthetic datasets above we do not apply this strategy, i.e., $n=1$, as the sample size is relatively smaller. The mini-batch size is set to 256 for all experiments, except for Dyngen, where we use a smaller batch size of 60 due to its limited number of samples.
Comparison with sinkhorn divergence as distribution fitting loss $\mathcal{L}_{\rm OT}$. We compute the distribution fitting loss $\mathcal{W}_{1}$ in Sect. 3.4 using EMD distance by pot library, following MIOFlow [huguet2022manifold] and DeepRUOT [zhang2024learning]. Particularly, we compute the OT using the pot library, using the function pot.emd(). This function solves for $\pi$ using a network flow algorithm, which is not compatible with PyTorch’s automatic differentiation. However, the gradients of the cost function $c$ can still be backpropagated.
We replace the EMD distance with Sinkhorn divergence [feydy19a], which is fully differentiable and enjoys many favorable properties, using CUDA/C++ based geomloss library. Tab. A-10 verified the effectiveness of Sinkhorn divergence for the distribution fitting loss.
Sinkhorn divergence between distributions $p_{0}$ and $p_{1}$ is defined as $S_{\epsilon}(p_{0},p_{1})=\mathcal{W}_{\epsilon}(p_{0},p_{1})-\frac{1}{2}\mathcal{W}_{\epsilon}(p_{0},p_{1})-\frac{1}{2}\mathcal{W}_{\epsilon}(p_{1},p_{1})$, where $\mathcal{W}_{\epsilon}(p_{0},p_{1})$ is the entropy-regularized optimal transport distance (Sinkhorn distance) and $\epsilon>0$ is the entropy regularization parameter.
Table A-10: Comparison of EMD and sinkhorn divergence (with different entropic regularization parameters) on EB 50D dataset with metrics ($\mathcal{W_{1}}$/RME at each timepoint).
$\mathcal{L}_{\rm OT}$ $t_{1}$ $t_{2}$ $t_{3}$ $t_{4}$ Training time(min)
EMD 7.951/0.039 8.747/0.042 9.244/0.019 9.620/0.044 13
$S_{\epsilon}(\epsilon=0.001)$ 7.902/0.018 8.767/0.013 9.063/0.083 9.507/0.096 9
$S_{\epsilon}(\epsilon=0.002)$ 7.904/0.032 8.791/0.033 9.111/0.104 9.523/0.120 9
$S_{\epsilon}(\epsilon=0.005)$ 7.917/0.048 8.768/0.062 9.086/0.137 9.518/0.151 9
$S_{\epsilon}(\epsilon=0.01)$ 7.904/0.035 8.801/0.038 9.102/0.113 9.511/0.127 9
$S_{\epsilon}(\epsilon=0.05)$ 7.908/0.036 8.787/0.045 9.080/0.120 9.512/0.129 9
$S_{\epsilon}(\epsilon=0.1)$ 7.916/0.030 8.773/0.022 9.111/0.091 9.591/0.108 9
It can be found that the sinkhorn divergence achieves further improvement on the accuracy of dynamics reconstruction, and costs less time for computing the distribution fitting loss.
(a) Predicted dynamics and predicted growth rates by VGFM on 5D EB data with timepoint 1 held out.
@@
We also examine whether the distribution fitting loss might lead to overfitting on the given population snapshots. In the hold-one-out experiments (Tab. 2), we relabel the timestamps of the four timepoints as 0, 1, 2, 3, and 4. Following the protocol in [kapusniak2024metric], we evaluate the model by holding out one intermediate timepoint at a time (i.e., 1, 2, and 3) and report the average Wasserstein-1 distance. While hold-one-out experiments have already validated the superiority of VFGM over other methods, we further analyze the loss curves of both the distribution fitting loss and the $\mathcal{W}_{1}$ distance on a hold-out distribution (e.g., the distribution at time 1) that is not seen during training (Fig. A-18). The results show that the $\mathcal{W}_{1}$ distance on the hold-out distribution decreases during the optimization of $\mathcal{L}_{\rm OT}$, indicating the absence of overfitting and demonstrating the generalization ability of the distribution fitting loss in modeling unseen distributions.
@@ B.7 CITE Data
CITE-seq (Cellular Indexing of Transcriptomes and Epitopes by Sequencing) [lance2022multimodal] is an advanced technique that enables the simultaneous profiling of transcriptomes and surface protein expression at the single-cell level through the use of antibody-derived tags. In this study, we utilize only the gene expression matrix from the CITE-seq dataset, and preprocess the data by normalizing CITE (5D) and leaving CITE (50D) unnormalized. The experimental settings for the CITE dataset are consistent with those used in the EB data experiments, employing PCA with 5 and 50 dimensions, and hyperparameters set as $\epsilon=0.01$, $\tau=5$. We relabel the timestamps of the four timepoints as 0, 1, 2, and 3. Following the protocol in [kapusniak2024metric], we evaluate the model by holding out one intermediate timepoint at a time (i.e., 1 and 2) and report the average Wasserstein-1 distance.
@@
To assess the model’s capability to capture state trajectories (Fig. 20(a) (a)) and predict mass growth (Fig. 20(a) (b)), the distribution at time point 1 is held out. Additionally, the distribution at time point 2 is held out to evaluate the performance of the $\mathcal{L}_{\rm OT}$ (Fig. 20(b)). The strong performance observed in the visualizations substantiates the potential of the proposed VGFM framework.
(a) Predicted dynamics and predicted growth rates by VGFM on 5D CITE Data with timepoint 1 held out.
(b) Visualization of predicted dynamics by (a) VFGM (w/o $\mathcal{L}_{\rm OT}$) and (b) VGFM on CITE (50D) dataset, where the hold-out time is the second intermediate timepoint.
B.8 Pancreas Data Analysis
To further explore the scalability of VGFM, we applied our model and compared methods explicitly modeling $g_{t}(x)$ [sha2024reconstructing, zhang2024learning] to the Pancreas dataset [bastidas2019comprehensive] comprising measurements of the developing mouse pancreas. We select cells at day 14.5 and 15.5 and relabel them as 0, 1 and select 2000 highly variable genes. As shown in [eyringunbalancedness], we assume the following cell type transitions are exclusively correct (denoted by $\rightarrow$), i.e., there is no descending cell type (or set of cell types) other than the given one. We partition all considered cell type transitions into three branches.
Endocrine branch (ED) transitions.
  • Fev+ Alpha (FA) $\rightarrow$ Alpha (A)
  • Fev+ Beta (FB) $\rightarrow$ Beta (B)
  • Fev+ Delta (FD) $\rightarrow$ Delta (D)
  • Fev+ Epsilon (FE) $\rightarrow$ Epsilon (E)
  • A $\rightarrow$ A
  • B $\rightarrow$ B
  • D $\rightarrow$ D
  • E $\rightarrow$ E
Ngn3 EP transitions.
  • Ngn3 high early (NE) $\rightarrow$ ED
  • Ngn3 high late (NL) $\rightarrow$ ED
Non-endocrine branch (NEB) transitions.
  • Ductal (DU) $\rightarrow$ DU
  • Tip (T) $\rightarrow$ Acinar (AC)
  • AC $\rightarrow$ AC
We observed that VGFM is the only method showing a steadily decreasing training loss, both for $\mathcal{L}_{\mathrm{VGFM}}$ and $\mathcal{L}_{\mathrm{OT}}$. We report $\mathcal{W}_{1}$ and RME as shown in Tab. A-11.
Table A-11: $\mathcal{W}_{1}$ and RME at day 15.5 between real data and generated data from VGFM and its variant.
Method $\mathcal{W}_{1}$ RME
VGFM (w/o $\mathcal{L}_{\rm OT})$ 27.002 0.025
VGFM 24.416 0.017
This indicates that both loss terms are still effective in high-dimensional real-world data.
Analysis on mean and variance. We calculated the means and variances of the real and generated gene at day 15.5 and plotted the corresponding mean-variance trend (Fig. 24(a) (a), (b)) and histograms (Fig. 24(a) (c), (d)). The results show that the generated samples closely follow the mean–variance trend of the real data, especially by incorporating $\mathcal{L}_{\rm OT}.$
(a) mean-variance trend ((a), (b)) and histograms ((c), (d)) of real and generated gene data.
Interpretable learned growth function. We highlight that our main contribution lies in modeling and training the growth function $g_{t}(x)$, with VGFM in the flow matching framework, to leverage snapshot mass change for learning $g_{w}(x,t)$. During training, the growth loss rapidly converges to a negligible value, much faster than the velocity loss. To further investigate $g_{w}(x,t)$, We first calculate growth rate of each cell at the initial time point (day 14.5) and visualize them in Fig. A-24. At this time, the proliferation observed during this developmental stage mainly originates from Acinar and Ductal cells. Remarkably, our model successfully recapitulates this pattern without being given any prior knowledge of the cell types, demonstrating its ability to infer biologically meaningful dynamics directly from the data.
Figure A-24: Normalized growth rate of each cell at time point 0 (day 14.5).
We also demonstrate that the learned $g_{w}(x,t)$ not only successfully recapitulates the relative growth rates of different cell types without any prior information about cell identities, but also that the branch-wise mass obtained from numerical integration of the ODE closely matches the total branch mass at time point 1.
Figure A-25: Relative mass of different branches
we then follow [sha2024reconstructing], for each time $t$, compute the element-wise absolute value $\left|\frac{\partial g(x,t)}{\partial x_{i}}\right|$ in the gene space for every cell, and average across cells to quantify each gene’s contribution to growth dynamics. From our analysis of $g_{w}(x,t)$, We have identified Pnliprp1, Clps, and Ctrb1, as shown in Fig. 27(a). These genes are well-known markers of Acinar cells and are highly expressed in the exocrine pancreas. Notably, as shown above, cells from the non-endocrine branch, particularly Acinar and Ductal cells, are much more abundant at later stages due to their high proliferation rates. This alignment between the learned growth-driving genes and known biological processes indicates that VGFM captures growth dynamics in a biologically interpretable manner, successfully linking the latent growth function to meaningful cell-type–specific proliferation patterns.
(a) Key genes identified by $g_{w}(x,t)$ at time 0 (a) and 1 (b).